Man is five times older than his son. Four years ago he was nine times older than his son. Find thier present ages.
Answer by mananth(12270)

let son be x years
father =5x
five years ago their ages were
x-5 & 5x-5
5x-5=9(x-5)
5x-5=9x-45
4x=40
x=10 years ---- son's age
Father's age = 50 years